OpenStudy (anonymous):

A mini toy helicopter ascends 50 feet vertically, then flies horizontally 200 feet. Find the angle of elevation to the helicopter now as seen by an observer at the takeoff point. Round to the nearest degree

OpenStudy (amistre64):

one of the trigs uses up and over in its definition .... woulnt the inverse of it tell us the angle?

OpenStudy (anonymous):

\[\frac{180 \tan ^{-1}\left(\frac{50}{200}\right)}{\pi }=14. {}^{\circ} \]